# CardioLab: Laboratory Values Estimation from Electrocardiogram Features - An Exploratory Study

### v1

@@ 1 Introduction
Abnormal laboratory values play a crucial role in clinical settings as they indicate underlying health conditions, sometimes posing severe and potentially life-threatening risks, especially in cases of serious illnesses [1, 2]. These abnormal values significantly increase the risk of mortality compared to normal ranges [3], highlighting the need for vigilant monitoring and effective management strategies. Continuous monitoring of patients often involves inherent limitations such as invasive procedures like frequent blood sampling, which are resource-intensive [4] and suffer from delays between sampling and obtaining results [5] or potentially missing staffing of laboratory medicine during the night, thereby limiting real-time monitoring capabilities.
While previous research has explored the correlation between abnormal laboratory values and changes in electrocardiogram (ECG) readings [6, 7], the full potential of utilizing ECG data for accurately estimating these abnormalities remains largely untapped. Therefore, in this work, we propose a first exploratory study in which we investigate the feasibility of estimating laboratory values from ECG features with the addition of other data modalities such as patient demographics.
2 Background
@@ 3 Methods
Dataset The dataset comprises data sourced from the publicly available MIMIC-IV [18, 19] and MIMIC-IV-ECG [20, 19] datasets. It includes a comprehensive set of non-invasive features: demographics such as gender, age, and race (Caucasian, African, Asian, Latino, or other); ECG features including RR interval, P onset, P end, QRS onset, QRS end, and T end (all in milliseconds), as well as P axis, QRS axis, and T axis (in degrees). In terms of target abnormalities, we work on binary classification cases where we define a positive case when the considered laboratory value is lower or higher than the median low or high threshold values (provided as part of MIMIC-IV) across all samples of the same laboratory value. For dataset splitting, we follow the stratification based on patients, age, and gender from the previously proposed MIMIC-IV-ECG-ICD [21] splits with a train, validation, and test ratio of 18:1:1 which ensures balanced diagnoses in each of the folds. Finally, for this work, we consider final laboratory value cases where we obtain at least 10 positive and 10 negative cases per fold.
Models and performance evaluation We fit and train individual extreme gradient boosting (XGBoost) tree models with a max depth of 1 per laboratory value in a binary classification setting. We evaluate performances on the macro average across all areas under the respective receiver operating curves (AUROC) (macro AUROC), where to assess statistical uncertainty resulting from the finite size and specific composition of the test set, we employ empirical bootstrap on the test set with $n=1000$ iterations and report 95% confidence intervals.
@@ 4 Results
Table 1: Performance results for laboratory values with AUROC > 0.70, presented by value name, abnormality setting, threshold defining abnormality by the number of units, (units), number of samples, class prevalence, and AUROC with 95% confidence intervals. All of these values represent blood fluid.
Value Threshold Unit Samples [Prev.] AUROC (95% CI)
Urea Nitrogen <6.0 mg/dL 240892 [2139] 0.856 (0.855, 0.857)
NTproBNP >353.0 pg/mL 23206 [18035] 0.848 (0.848, 0.849)
Hematocrit >51.0 % 238567 [1381] 0.775 (0.773, 0.777)
Cortisol >20.0 ug/dL 982 [486] 0.771 (0.772, 0.781)
Osmolality <275.0 mOsm/kg 2784 [730] 0.765 (0.763, 0.768)
Hemoglobin >17.5 g/dL 236802 [802] 0.754 (0.753, 0.759)
Creatinine <0.5 mg/dL 241968 [2499] 0.742 (0.741, 0.744)
Urea Nitrogen >20.0 mg/dL 240892 [100216] 0.736 (0.736, 0.736)
Creatinine >1.2 mg/dL 241968 [69012] 0.735 (0.735, 0.736)
PT <9.4 sec 137917 [380] 0.732 (0.728, 0.734)
C-Reactive Protein >5.0 mg/L 3520 [2107] 0.730 (0.728, 0.734)
Anion Gap >20.0 mEq/L 230666 [18299] 0.726 (0.725, 0.727)
Cholesterol, HDL <41.0 mg/dL 9023 [2103] 0.725 (0.725, 0.729)
Lymphocytes >42.0 % 171441 [9591] 0.722 (0.721, 0.723)
RDW-SD >46.3 fL 73001 [35887] 0.720 (0.719, 0.721)
Bands >5.0 % 16197 [2587] 0.719 (0.718, 0.721)
Bicarbonate, Whole Blood <21.0 mEq/L 8039 [1720] 0.715 (0.712, 0.717)
Ferritin <13.0 ng/mL 4643 [438] 0.714 (0.714, 0.722)
Lymphocytes <18.0 % 171441 [73323] 0.713 (0.712, 0.713)
Monocytes >11.0 % 9898/171442 0.713 (0.712, 0.714)
Bilirubin, Direct > 0.3 mg/dL 3438 [1968] 0.708 (0.704, 0.710)
Albumin <3.5 g/dL 57040 [12834] 0.708 (0.707, 0.708)
Sodium >145.0 mEq/L 235578 [5361] 0.706 (0.705, 0.707)
eAG >123.0 mg/dL 9485 [4916] 0.704 (0.702, 0.705)
Ferritin >150.0 ng/mL 4643 [2013] 0.704 (0.701, 0.705)
% Hemoglobin A1c >5.9 % 10665 [5651] 0.700 (0.699, 0.701)
Overall Predictive Performance Table 1 contains the performance results in terms of AUROC for individual laboratory value abnormalities above 0.70. This includes the threshold for their label definition, units, number of samples, and prevalence. Overall, low urea nitrogen and NTproBNP, achieve an AUROC above 0.80. The table lists 26 laboratory values with AUROC greater than 0.70, indicating strong predictive performance. These values span various physiological systems, demonstrating the model’s capability to predict abnormalities across diverse domains beyond cardiovascular (NTproBNP, hematocrit, hemoglobin), coagulation (PT), and electrolyte balance (osmolality, anion gap, sodium). Notably, the model also encompasses renal (urea nitrogen, creatinine), endocrine (cortisol, eAG), immune (bands, monocytes, lymphocytes), inflammatory response (C-Reactive protein), metabolic (cholesterol, bicarbonate), and respiratory (pO2) systems. This broad applicability underscores the model’s extensive utility in clinical prediction across a wide range of physiological parameters.
Low urea nitrogen use-case: predictive performance for different patient demographics By gender, males perform better than females despite having a smaller prevalence rate. By race, the groups of Caucasians and others perform above the overall population mean, Caucasians however, with the largest prevalence, Asians have the lowest predictive performance, however, also the ones with the fewest samples, to make a fair comparison, we down-sampled the Caucasian training population to Asians population, and conclude that indeed Caucasian perform better than (at least) Asians, even with the same number of samples. By age, the 65 to 77 years age group is the most predictive across all groups despite having a small prevalence.
Low urea nitrogen use-case: interpretability XGBoost enables explainable methods like Shapley additive explanations (SHAP) [22]. Using SHAP, we identify ECG features contributing most to abnormality predictions. Literature indicates hemodialysis patients with reduced urea nitrogen levels often show ECG abnormalities [23]. Our findings suggest: (1) High T end values (long T wave) correlate with prolonged repolarization abnormalities, potentially linked to observed ST segment changes (depression or elevation) and inverted T waves in these patients. (2) Low T axis values correlate with T wave inversions and other repolarization abnormalities. (3) High QRS axis values indicate a shift in ventricular depolarization’s mean electrical axis, possibly associated with ventricular hypertrophy or conduction abnormalities in hemodialysis patients. Furthermore, high NTproBNP values, our second-best performer, align with literature associating abnormal ECG features in a specific cohort [24].
@@ 5 Discussion
Clinical significance This work enhances ECG predictive capabilities beyond traditional applications, laying the groundwork for comprehensive patient monitoring as in [25] for diagnostics. By analyzing ECG data alongside patient demographics, we predict abnormalities in crucial laboratory markers such as electrolyte levels (e.g., Anion gap, Sodium, Bicarbonate) and renal function indicators like urea nitrogen and creatinine. This predictive approach facilitates early detection of conditions such as arrhythmias due to electrolyte imbalances or kidney dysfunction, enabling proactive interventions that can optimize patient care by providing faster and low-cost results [26].
Limitations We identify several limitations in our current work, firstly, for clinical consideration, one might need to externally validate these models on a separate population cohort. Secondly, we aim to estimate the closest absolute value within a horizon threshold, however, fixed intervals between features and targets for all samples might mitigate variance. Lastly, A more fine-grained setting can include drug-specific pharmacokinetic information to account for drug effects in the target space. However, the complexity of it would increase significantly as half-life, and mechanisms of action and elimination are drug-specific.
Future work While this work investigates the predictive performance of laboratory values by discriminating ECG features within a stratified population mostly presenting diagnoses [21], it is also important to understand how changes in ECG features contribute to predictions in a population of only healthy individuals [27]. Additionally, previous research demonstrates the superiority of using ECG signals over features for diagnosis and decompensation prediction [28]. Extending this work to include waveforms rather than relying solely on tabular data would be a valuable direction for future research.
@@
Appendix
Table 2: Performance results by demographic groups, presented by category, the number of samples, prevalence, and AUROC with 95% confidence intervals.
Category Samples [Prev.] AUROC
Gender
Males 122,125 [686] 0.844 (0.842, 0.846)
Females 118,767 [1453] 0.829 (0.828, 0.831)
Race
Caucasians 157,926 [1287] 0.881 (0.880, 0.882)
Africans 40,205 [432] 0.784 (0.784, 0.790)
Asians 7,295 [63] 0.653 (0.652, 0.666)
Latinos 13,842 [149] 0.763 (0.765, 0.774)
Other 17,306 [206] 0.871 (0.869, 0.873)
Age by quantiles
(18-49yo.) 54,830 [1181] 0.723 (0.720, 0.723)
(50-64yo.) 54,203 [617] 0.702 (0.701, 0.708)
(65-77yo.) 65,854 [227] 0.808 (0.805, 0.815)
(>=78yo.) 52,596 [114] 0.709, (0.705, 0.708)

### v2

@@ 1 Introduction
Clinical relevance Abnormal laboratory values play a crucial role in clinical settings as they indicate underlying health conditions, sometimes posing severe and potentially life-threatening risks, especially in cases of serious illnesses [1, 2]. These abnormal values significantly increase the risk of mortality compared to normal ranges [3], highlighting the need for vigilant monitoring and effective management strategies. Continuous monitoring of patients often involves inherent limitations such as invasive procedures like frequent blood sampling, which are resource-intensive [4] and suffer from delays between sampling and obtaining results [5] or potentially missing staffing of laboratory medicine during the night, thereby limiting real-time monitoring capabilities.
Table 1: Performance results for the top 25 laboratory values presented by value name, abnormality setting, threshold defining abnormality by the number of units, (units), number of samples, class prevalence, and AUROC with 95% confidence intervals.
Value Threshold Unit Samples [Prev.] AUROC (95% CI)
Albumin >5.2 g/dL 57040[0.4%] 0.901 (0.901, 0.906)
Hemoglobin >17.5 g/dL 236802[0.34%] 0.881 (0.881, 0.884)
NTproBNP >353.0 pg/mL 23206[77.72%] 0.866 (0.865, 0.867)
Acetaminophen >30.0 ug/mL 1205[18.76%] 0.857 (0.856, 0.864)
Hematocrit >51.0 % 238567[0.58%] 0.847 (0.846, 0.849)
PT <9.4 sec 137917[0.28%] 0.831 (0.828, 0.831)
Red Blood Cells >6.1 m/uL 234079[0.31%] 0.818 (0.818, 0.822)
25-OH Vitamin D <30.0 ng/mL 1101[55.4%] 0.815 (0.813, 0.822)
RDW-SD <35.1 fL 73001[0.22%] 0.811 (0.811, 0.817)
INR(PT) <0.9 nan 137973[0.22%] 0.803 (0.796, 0.801)
Urea Nitrogen <6.0 mg/dL 240892[0.89%] 0.796 (0.796, 0.798)
Monocytes >11.0 % 171442[5.77%] 0.768 (0.768, 0.769)
Acetaminophen <10.0 ug/mL 1205[48.22%] 0.768 (0.768, 0.775)
Absolute Basophil Count <0.01 K/uL 51315[7.81%] 0.765 (0.765, 0.767)
Urea Nitrogen >20.0 mg/dL 240892[41.6%] 0.756 (0.756, 0.756)
C-Reactive Protein >5.0 mg/L 3520[59.86%] 0.753 (0.752, 0.757)
Cholesterol, HDL <41.0 mg/dL 9023[23.31%] 0.749 (0.746, 0.749)
Bilirubin, Direct >0.3 mg/dL 3438[57.24%] 0.748 (0.747, 0.752)
RDW-SD >46.3 fL 73001[49.16%] 0.744 (0.744, 0.745)
Hemoglobin <13.7 g/dL 236802[72.39%] 0.741 (0.741, 0.742)
Creatinine >1.2 mg/dL 241968[28.52%] 0.738 (0.738, 0.739)
Sedimentation Rate >20.0 mm/hr 1861[57.87%] 0.736 (0.734, 0.74)
pO2 <85.0 mm Hg 35047[46.75%] 0.733 (0.732, 0.733)
Osmolality, Measured <275.0 mOsm/kg 2784[26.22%] 0.729 (0.726, 0.732)
Bicarbonate >32.0 mEq/L 231772[2.86%] 0.728 (0.728, 0.729)
@@
Contribution While previous research has explored the correlation between abnormal laboratory values and changes in electrocardiogram (ECG) readings [16, 17], the full potential of utilizing ECG data for accurately estimating these abnormalities remains largely untapped. Therefore, in this work, we propose a first exploratory study in which we investigate the feasibility of estimating laboratory values from ECG features with the addition of other data modalities such as patient demographics.
@@ 2 Methods
Dataset The dataset comprises data sourced from the publicly available MIMIC-IV [18, 19] and MIMIC-IV-ECG [20, 19] datasets. It includes a comprehensive set of non-invasive features: demographics such as gender, age, and race, as well as vital signs such as temperature, heart rate, respiration rate, oxygen saturation, diastolic blood pressure, systolic blood pressure; and finally, ECG features including RR interval, P onset, P end, QRS onset, QRS end, and T end (all in milliseconds), as well as P axis, QRS axis, and T axis (in degrees). In terms of target abnormalities, we work on binary classification cases where we define a positive case when the considered laboratory value is lower or higher than the patient-wise median low or high threshold values (provided within MIMIC-IV) across all samples of the same laboratory value. For data sampling, the estimation task involves using the nearest vital signs recorded within a 30-minute interval from the ECG data to predict laboratory abnormalities. These predictions are based on the closest lab values within a 60-minute window. For dataset splitting, we follow the patient-based stratification based on demographics and diagnoses proposed by MIMIC-IV-ECG-ICD [21], which splits with a train, validation, and test ratio of 18:1:1. Finally, for this work, we consider final laboratory value cases where we obtain at least 10 positive and 10 negative cases per fold.
Models and performance evaluation We fit and train individual extreme gradient boosting (XGBoost) tree models with a max depth of 1 per laboratory value in a binary classification setting. We evaluate performances based on areas under the respective receiver operating curves (AUROC). To assess statistical uncertainty resulting from the finite size and specific composition of the test set, we use empirical bootstrapping on the test set with $n=1000$ iterations and report 95% confidence intervals.
@@ 3 Results
Overall Predictive Performance Table 1 contains the performance results in terms of AUROC for the 25 best-performing individual laboratory value abnormalities. This includes the threshold for their label definition, units, number of samples, and prevalence. This selection of laboratory values, all of which represent blood fluids, underscores the model’s capacity to predict abnormalities in diverse bodily systems. Notable among these are values related to cardiovascular function (Albumin, Hemoglobin, NTproBNP), coagulation (PT, INR), and oxygen transport (pO2). The model also effectively addresses renal function (Urea Nitrogen, Creatinine), immune response (Monocytes, Absolute Basophil Count), and inflammatory processes (C-Reactive Protein, Sedimentation Rate). Moreover, it accurately identifies metabolic (Cholesterol, HDL, Bicarbonate, Osmolality), endocrine (25-OH Vitamin D), and hepatic (Bilirubin, Direct) abnormalities. The inclusion of values like Acetaminophen and RDW-SD highlights the model’s versatility in managing both common and more specialized laboratory measurements.
@@ 4 Discussion
Clinical significance This work enhances ECG predictive capabilities beyond traditional applications, in line with its counterpart [22] that demonstrates the predictability in particular of non-cardiac conditions from ECG features and clinical metadata. The model’s ability to predict a wide range of clinically significant laboratory abnormalities using only ECG features, demographics, and basic, non-invasive vital signs is a major advancement in healthcare. By identifying critical cardiovascular issues such as those related to albumin and hemoglobin levels, the model enables early detection and intervention without the need for expensive or invasive testing. Its capacity to predict heart failure risk through NTproBNP levels and identify coagulation abnormalities, such as those indicated by PT and INR, further demonstrates its potential to improve patient outcomes by facilitating timely and targeted care. Moreover, the model’s ability to assess renal function, immune response, and inflammatory processes using easily obtainable data could significantly enhance diagnostic accuracy and accessibility, particularly in resource-limited settings. This approach not only streamlines clinical workflows but also broadens the scope of preventive care, making it easier to monitor and manage patient’s health effectively while providing faster and low-cost results.
Limitations We identify several limitations in our current work. Firstly, for clinical consideration, the present promising results will have to be externally validated based on a separate population cohort. Secondly, we estimate the lab value with the closest time difference to the point in time where the ECG was taken. Using a fixed intervals between features and targets for all samples might mitigate variance. Thirdly, our work builds on constant lower and higher abnormality thresholds extracted from MIMIC. More reliable threshold choices, also depending on patient characteristics such as age and gender, should be investigated in the future.
Future work
Firstly, our proposed method enables an easy integration of explainable methods such as Shapley additive explanations (SHAP) [23] which will allow us to identify the specific set of features that contributing most to abnormality predictions. This would allow to quantify the relative impact of the different input modalities, demographics, vital signs, or ECG features on the prediction performance.
Secondly, in addition to the assessment of the overall model performance, it will be very instructive to also investigate the performance on specific subgroups defined for example based on demographics such as gender, ethnicity, and age groups. However, one must consider that certain groups might suffer from biased results such as abnormal values being considered normal at certain aging stages, which goes in line with more finegrained threshold definitions as discussed above.
Thirdly, while this work investigates the predictive performance of laboratory values by discriminating ECG features within a largely critically ill, it might be informative to investigate the same question in a healthy subgroup. This parallels recent efforts to characterize age-related ECG changes in a healthy cohort [24].
Finally, previous research has demonstrated the superiority of using ECG waveforms over ECG features for diagnosis and decompensation prediction [25]. Extending this work to include waveforms rather than relying solely on tabular data would be a valuable direction for future research.